OpenStudy (anonymous):

Slope question...how do I do this D:

OpenStudy (anonymous):

Write an equation of the line that passes through the point (-4,-1) and is (a) parallel to and (b) perpendicular to the line 2x+7y=14

OpenStudy (anonymous):

so...what do I do?

OpenStudy (anonymous):

At first you consider the conditions a and b. From a you know that the equation that you are looking for has the same format as the one you gave in the exercise, because the slope must be the same, but with an addition to make the line go up or down as you need. In a you only need to find that addition. Do you agree?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Let me reread that..

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Parallel has the same slope...so...

OpenStudy (anonymous):

Do i change it to....slope intercept form?

OpenStudy (anonymous):

You can, its up to you, it doesn't really change anything mathematically speaking.

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

so y=-2/7x+2?

OpenStudy (anonymous):

And then use the coordinates to make point slope form and then perpendicular?

OpenStudy (anonymous):

Well I didn't understood what you meant now, but you need the coordinates to find the new value for that 2 in the end of the formula. For perpendicular, you need to find the new slope.

OpenStudy (anonymous):

For perpendicular..don't i just do reciprocal?? ...

OpenStudy (anonymous):

is Parallel y=-2/7x+2 Perpendicular y=7/2x+13

OpenStudy (anonymous):

Yes, thats it, but you still need to change that 2 because this line don't pass throguh that point.

OpenStudy (anonymous):

Same thing with the second equation.